What is the general form of the equation of the given circle with center (-3,12) and the radius as 5.
Answer by nerdybill(7384)

What is the general form of the equation of the given circle with center (-3,12) and the radius as 5.
.
Standard orm of a circle is:
(x-h)^2 + (y-k)^2 = r^2
where
(h,k) is the center
and
r is the radius
.
(x-h)^2 + (y-k)^2 = r^2
plug in the given center (-3,12) and radius 5
(x-(-3))^2 + (y-12)^2 = 5^2
(x+3)^2 + (y-12)^2 = 25 (general form)
.
expand to get "general form"
x^2 + y^2 + Dx + Ey + F = 0
(x+3)^2 + (y-12)^2 = 25
(x+3)(x+3) + (y-12)(y-12) = 25
x^2+6x+9 + y^2-24y+144 = 25
x^2+y^2+6x-24y+153 = 25
x^2+y^2+6x-24y+128 = 0 (answer)